Anna holds a 13 D magnifier directly in front of her eye to get a close look at a 19-mm -diameter penny. Assume a typical 25 cm near point and a distance of 1.7 cm between the lens and the retina.

Part A

What is the closest possible distance that she can hold the coin to have it appear in focus?

Express your answer with the appropriate units.

Please answer and show full work/explain why! Thanks :)
Answer #1

P = Power = 13 D

f = focal length = 100/P = 100/13 = 7.69 cm

di = image distance = - 25 - 1.7 = - 26.7 cm

do = object distance = ?

Using the equation

1/f = 1/di + 1/do

1/7.69 = 1/(- 26.7) + 1/do

do = 5.97 cm